# Variational Control for Guidance in Diffusion Models

### v1

@@ 4.1 Non-linear Inverse Problems
Gaussian blur Motion blur
Method LPIPS↓ FID↓ Time↓ LPIPS↓ FID↓ Time↓
DMPlug 0.147 69.36 51.24 0.118 72.85 51.13
NDTM (Ours) 0.103 55.15 7.17 0.086 49.99 7.17
NDTMζ (Ours) 0.083 47.34 18.07 0.063 38.6 18.13
Blind Image Deblurring (BID). Next, we extend our framework to blind image deblurring, maintaining the same setup as DMPlug (Wang et al., 2024), which is the state-of-the-art method for this task. Interestingly, adapting our proposed method for blind inverse problems involves jointly optimizing the unknown blur kernel parameters along with the control ${\mathbf{u}}_{t}$. More specifically, for degradation of the form ${\bm{y}}={\bm{k}}*{\mathbf{x}}_{0}+\sigma_{y}{\mathbf{z}}$ with unknown blurring kernel ${\bm{k}}$, we update line 11 in Algorithm 1 as ${\mathcal{C}}_{\text{terminal}}=w_{T}\Phi(\hat{{\mathbf{x}}}_{0}^{(i)},k^{(i)})$ and optimizing for the trainable kernel for each image, as ${\bm{k}}^{(i+1)}\leftarrow\texttt{Update}({\bm{k}}^{(i)},\nabla_{k}{\mathcal{C}}_{t}^{(i)}$).
Figure 2 (Right) illustrates the comparison between DMPlug and NDTM adapted for this task. Qualitatively, we find that while DMPlug can introduce artifacts in generating reconstructions, NDTM generates high-quality reconstructions. Table 2 further validates our qualitative findings as our method outperforms DMPlug on perceptual quality metrics. More interestingly, while DMPlug is extremely expensive for a single image, our method outperforms the former on sample quality by a significant margin while being an order of magnitude faster. This illustrates that our sampler has a more efficient way to trade sampling speed for quality. We additionally present quantitative results for a more efficient configuration ($N=15,T=200$) of our proposed method that maintains strong performance in Table 2, with corresponding qualitative results provided in Appendix C.
Figure 2: NDTM outperforms competing baselines on noisy blind image deblurring (BID) with Gaussian (top) and Motion (bottom) kernels. NDTM accurately captures most details, while competing methods introduce artifacts in the generated reconstructions.
@@
4.2 Linear Inverse Problems
Lastly, we compare competing methods on linear inverse problems: (4x) Super-resolution and Random inpainting with a 90% masking probability. As illustrated in Table 3, for super-resolution, NDTM outperforms competing baselines for both datasets. For random inpainting, our method performs comparably with DPS on the FFHQ dataset. However, for a more difficult benchmark like ImageNet, NDTM outperforms the next best-competing baseline, DPS on this task. We present additional qualitative results for linear inverse problems in Appendix C
4.3 Ablation Studies
@@
Next, we analyze the impact of different design choices in NDTM on the perception (LPIPS) and distortion (PSNR) quality for the non-linear deblur task on the ImageNet dataset.
Impact of Guidance. Since the terminal cost weight $w_{T}$ and the parameter $\gamma$ affect the optimization of the variational control parameters ${\mathbf{u}}_{t}$, we analyze their impact on sample quality. From Fig. 3a, we observe that increasing the terminal weight $w_{T}$ leads to an improvement in both perceptual and distortion quality. However, in the limit of $w_{T}\rightarrow\infty$ (i.e., where the regularization terms in Eq. 17 can be ignored), the perceptual quality degrades, which highlights the importance of the transient cost in our framework. Similarly, increasing $\gamma$ also leads to an improved sample quality. However, a large $\gamma$ can also lead to overshooting which can lead to a degradation in sample quality.
@@
5 Related Work
Conditional Diffusion Models. In general, the conditional score $\nabla_{{\mathbf{x}}_{t}}\log p({\mathbf{x}}_{t}|{\bm{y}})$ needed for guided sampling can be learned during training (Saharia et al., 2022; Podell et al., 2024; Rombach et al., 2022) or approximated during inference. Here, we focus on training-free guidance during inference.
In this context, there has been some recent progress in approximating the noisy likelihood score (see Eq. 5) by approximating the diffusion posterior $p({\mathbf{x}}_{0}|{\mathbf{x}}_{t})$. For instance, DPS (Chung et al., 2022a) approximates the diffusion posterior by a Dirac distribution centered on Tweedie’s estimate (Efron, 2011). This has the advantage that the guidance can be adapted to linear and non-linear tasks alike. However, due to a crude approximation, DPS converges very slowly and, in our observation, could be unstable for certain tasks (see Table 1). Consequently, some recent work (Yu et al., 2023; Bansal et al., 2024) adds a correction term on top of the DPS update rule to better satisfy the constraints. In addition to being more principled, our proposed method instead directly estimates the guided posterior at each sampling step, thus sidestepping the limitations of DPS in the first place.
More recent work (Song et al., 2022b; Pandey et al., 2024b; Pokle et al., 2024; Boys et al., 2023) relies on expressive approximations of the diffusion posterior. While this can result in accurate guidance and faster sampling, a large proportion of these methods are limited to linear inverse problems. In contrast, our method can be adapted to generic inverse problems with some additional computational overhead. Lastly, another line of work in inverse problems approximates the data posterior $p({\mathbf{x}}_{0}|{\mathbf{y}})$ using variational inference (Zhang et al., 2018a). For instance, RED-diff (Mardani et al., 2023a) proposes to learn an unimodal approximation to the data posterior by leveraging a diffusion prior. However, this can be too restrictive in practice and comes at the expense of blurry samples. We refer interested readers to Daras et al. (2024) for a more detailed review of training-free methods for solving inverse problems in diffusion models.
Optimal Control for Diffusion Models. There has been some recent interest in exploring connections between stochastic optimal control and diffusion models (Berner et al., 2024). Chen et al. (2024) leverage ideas from control theory for designing efficient diffusion models with straight-line trajectories in augmented spaces. Since our guided sampler can be used with any pretrained diffusion models, our approach is complementary to this line of work. More recently, SCG (Huang et al., 2024) leverages ideas from path integral control to design guidance schemes with non-differentiable constraints. In contrast, we only focus on differentiable terminal costs, and extending our framework to non-differentiable costs could be an important direction for future work. Lastly, Rout et al. (2024) propose RB-Modulation, a method based on control theory for personalization using diffusion models. Interestingly, while RB-Modulation is primarily inspired by a class of tractable problems in control theory, it is a special case of our framework in the limit of $w\rightarrow\infty$ and $\gamma=1$. Therefore, our proposed framework is more flexible.
@@ 6 Conclusion
Our novel framework elucidates the design space of guidance in diffusion models. Through the lens of variational optimal control, it offers a principled way to guide a pretrained diffusion model while respecting an external cost. Our empirical results suggest that optimizing each diffusion step allows for more flexibility in guidance compared to commonly used approximations of the diffusion posterior.
While our method adapts well to linear and (blind) non-linear problems, there remain several interesting directions for future work. First, a more thorough theoretical investigation into the optimization dynamics of the proposed method could help alleviate the extra overhead of variational inference during each diffusion sampling. Second, it would be interesting to extend our method to Latent Diffusion Models (LDMs) (Rombach et al., 2022). Lastly, our method is only one instantiation of our framework, leaving ample room for exploration. Investigating variational parameterizations, refining cost functions, and developing more efficient optimization schemes could further enhance the flexibility and applicability of our approach.
Acknowledgment. We thank Justus Will for additional discussions and feedback. This project was funded through support from the Chan Zuckerberg Initiative. Additionally, Stephan Mandt acknowledges funding from the National Science Foundation (NSF) through an NSF CAREER Award IIS-2047418, IIS-2007719, the NSF LEAP Center, the IARPA WRIVA program, and the Hasso Plattner Research Center at UCI.
@@ A.1 Simplification of the NDTM Objective for DDIM
In the case of DDIM (Song et al., 2022a), the diffusion posterior is parameterized as (Eqn. 12 in Song et al. (2022a)),
@@ A.2 Continuous-Time Diffusion Trajectory Matching
By (Song et al., 2021, Theorem 1 in Appendix A) (which follows from an application of Girsanov’s Theorem), the transient cost reads:
@@ B.1 Task Details
B.2 Task Specific Hyperparameters
Here, we provide a detailed overview of different hyperparameters for the baselines considered in this work. We optimize all baselines and our method for the best sample perceptual quality. We use the official code implementation for RED-Diff (Mardani et al., 2023a) at https://github.com/NVlabs/RED-diff, https://github.com/mandt-lab/c-pigdm and https://github.com/sun-umn/DMPlug for running all competing baselines.
@@ B.2 Task Specific Hyperparameters
We adopt the DPS parameters from Mardani et al. (2023a). More specifically, we fix the number of diffusion steps to 1000 using the DDIM sampler. We set $\eta=0.5$ for all tasks. Following Chung et al. (2022a), we set,
@@
Table 5 illustrates different hyperparams for DPS on all tasks for the FFHQ and ImageNet datasets.
@@
Following Kawar et al. (2022), we fix $\eta=0.85$, $\eta_{b}=1.0$, and the number of diffusion steps to 20 across all linear inverse problems.
@@
We set the number of diffusion steps to 20 for all tasks. It is also common to contract the reverse diffusion sampling for better sample quality by initializing the noisy state as proposed in Chung et al. (2022b). We denote the start time as $\tau$. We re-tune C-$\Pi$GDM for the best perceptual quality for all linear inverse problems. Table 5 illustrates different hyperparams for linear inverse problems. We find that C-$\Pi$GDM fails to recover plausible images for random inpainting task after numerous tuning attempts.
Table 4: DPS hyperparameters used for different tasks
@@
Table 5: C-$\Pi$GDM hyperparameters used for different tasks. We find that C-$\Pi$GDM fails to recover plausible images for this task after numerous tuning attempts.
@@
Task N $\gamma_{t}$ $\eta$ $\tau$ $w_{T}$ $w_{\text{score}}$ $w_{\text{control}}$
@@
B.2.4 RED-Diff (Mardani et al., 2023a)
We set $\sigma_{0}=0$ with a linear weighting schedule and $lr=0.5$, $\lambda=0.25$ and perform 50 diffusion steps for all tasks across the FFHQ and ImageNet dataset. We highlight different hyperparams in Table 6.
@@
Task N $\gamma$ $\eta$ $\tau$ $w_{T}$ $w_{\text{score}}$ $w_{\text{control}}$ N $\gamma$ $\eta$ $\tau$ $w_{T}$ $w_{\text{score}}$ $w_{\text{control}}$
@@
B.2.5 NDTM (Ours)
We use the Adam optimizer (Kingma & Ba, 2017) with default hyperparameters, fixing the learning rate to 0.01 for updating the control ${\mathbf{u}}_{t}$ across all tasks and fix the kernel learning rate in the BID task to 0.01. We refer to the loss weighting scheme in Eq. 16 as "DDIM weighting". Moreover, we use linear decay for the learning rate. We perform 50 diffusion steps across all datasets and tasks. We tune the guidance weight $\gamma$, the number of optimization steps N, loss weighting ($w_{T}$, $w_{\text{score}}$, $w_{\text{control}}$), DDIM $\eta$ and the truncation time $\tau$ (Chung et al., 2022b) for best performance across different tasks. All these hyperparameters are listed in Table 8.

### v2

@@ 5.1 Inverse Problems
Gaussian blur Motion blur
Method LPIPS↓ FID↓ Time↓ LPIPS↓ FID↓ Time↓
DMPlug 0.147 69.36 51.24 0.118 72.85 51.13
NDTM (Ours) 0.103 55.15 7.17 0.086 49.99 7.17
NDTMζ (Ours) 0.083 47.34 18.07 0.063 38.6 18.13
Non-Linear Deblurring. We consider non-linear deblurring with the same setup as in Chung et al. (2022a). For this task, we compare against DPS (Chung et al., 2022a), RED-Diff (Mardani et al., 2024), MPGD (He et al., 2024), and RB-Modulation (Rout et al., 2025) (which is a special case of our method with $\gamma_{t}=1.0$, $w_{s}=0$, $w_{c}=0$). Figure 1 (Right) illustrates the comparison between competing baselines and our proposed method, NDTM, on this task. Qualitatively, we find that DPS is very sensitive to guidance step size and is usually unstable on this task. Moreover, while RED-diff does not have stability issues, it is biased towards generating blurry samples. This is not surprising given their unimodal approximation to the data posterior $p({\mathbf{x}}_{0}|{\mathbf{y}})$. On the contrary, NDTM generates high-fidelity reconstructions with a stable sampling process. Similarly, our quantitative results in Table 1 validate our qualitative findings as our method outperforms competing baselines on perceptual quality for both datasets by a significant margin.
Blind Image Deblurring (BID). Next, we extend our framework to blind image deblurring, maintaining the same setup as DMPlug (Wang et al., 2024). We compare against DMPlug, which, to the best of our knowledge, is also the state-of-the-art method for this task. Interestingly, adapting our proposed method for blind inverse problems only involves jointly optimizing the unknown blur kernel parameters along with the control ${\mathbf{u}}_{t}$. More specifically, for degradation of the form ${\bm{y}}={\bm{k}}*{\mathbf{x}}_{0}+\sigma_{y}{\mathbf{z}}$ with unknown blurring kernel ${\bm{k}}$, we update line 11 in Algorithm 1 as ${\mathcal{C}}_{\text{terminal}}=w_{T}\Phi(\hat{{\mathbf{x}}}_{0}^{(i)},k^{(i)})$ and optimizing for the trainable kernel for each image, as ${\bm{k}}^{(i+1)}\leftarrow\texttt{Update}({\bm{k}}^{(i)},\nabla_{k}{\mathcal{C}}_{t}^{(i)}$).
Figure 2 (Right) illustrates the comparison between DMPlug and NDTM adapted for this task. Qualitatively, we find that while DMPlug can introduce artifacts in generating reconstructions, NDTM generates high-quality reconstructions. Table 3 further validates our qualitative findings as our method outperforms DMPlug on perceptual quality metrics. More interestingly, while DMPlug is extremely expensive for a single image, our method outperforms the former on sample quality by a significant margin while being an order of magnitude faster (see Table 3). This illustrates that our sampler has a more efficient way to trade sampling speed for quality. We present additional results in Appendix C.
Figure 2: NDTM outperforms competing baselines on blind image deblurring (BID) with Gaussian (top) and Motion (bottom) kernels. NDTM accurately captures most details, while competing methods introduce artifacts in the generated reconstructions.
Figure 3: Style Guidance with Stable Diffusion. NDTM (proposed) provides a better tradeoff between Prompt adherence and Style adherence using Stable Diffusion 1.4. While baselines can introduce random artifacts in generated samples, NDTM preserves sample quality while exhibiting better style adherence. (Top Panel) Reference Style Images. (Bottom Panel) Samples corresponding to the reference style images in the Top Panel. The CLIP (Radford et al., 2021) score (higher is better) represents the similarity between the generated image and the text prompt, indicating prompt adherence. The Style score (lower is better) represents the distance between CLIP features for the reference style and the generated image, indicating style adherence. We present additional quantitative results in Table 4.
Linear Inverse Problems. Lastly, we compare competing methods on linear inverse problems: (4x) Super-resolution and Random inpainting with a 90% masking probability. In addition to the baselines used for the non-linear deblur task, we also compare against DDRM (Kawar et al., 2022) and C-$\Pi$GDM (Pandey et al., 2024b). As illustrated in Table 2, for super-resolution, NDTM outperforms competing baselines for both datasets. For random inpainting, our method performs comparably with DPS on the FFHQ dataset. However, for a more difficult benchmark like ImageNet, NDTM outperforms the next best competing baseline, DPS, on this task. We present additional qualitative results for linear inverse problems in Appendix C
CLIP Score $\uparrow$ Style Score $\downarrow$
FreeDOM 30.86 508.28
MPGD 30.21 498.85
(Ours) NDTM 31.34 475.62
Table 4: Quantitative comparison between NDTM and other baselines on style guidance generation using Stable Diffusion 1.4. NDTM exhibits better prompt (see CLIP score) and style adherence (see Style Score) over competing baselines.
@@
5.2 Style Guidance
The goal in style guidance is to generate samples that adhere to a specified text prompt and the style of a reference image. We follow the same setup as MPGD and use a pretrained Stable Diffusion 1.4 text-to-image model. We specify the terminal cost for NDTM in Eq. 22 and compare with MPGD (He et al., 2024) and FreeDOM (Yu et al., 2023) on this task. Figures 3 and 4 demonstrate that NDTM exhibits a better balance between prompt and style adherence over competing baselines.
5.3 Ablation Studies
Next, we analyze the impact of different design choices in NDTM on the perception (LPIPS) and distortion (PSNR) quality for the non-linear deblur task on ImageNet.
Impact of Guidance. Since the terminal cost weight $w_{T}$ and the parameter $\gamma$ affect the optimization of the variational control parameters ${\mathbf{u}}_{t}$, we analyze their impact on sample quality. From Fig. 4a, we observe that increasing the terminal weight $w_{T}$ leads to an improvement in both perceptual and distortion quality. However, in the limit of $w_{T}\rightarrow\infty$ (i.e., where the regularization terms in Eq. 17 can be ignored), the perceptual quality degrades, which highlights the importance of the transient cost in our framework. Similarly, increasing $\gamma$ also leads to an improved sample quality. However, a large $\gamma$ can also lead to overshooting.
@@ 6 Conclusion
Our proposed framework offers a principled way to guide a pretrained diffusion model while respecting an external cost through the lens of variational optimal control. Our empirical results suggest that optimizing each diffusion step allows for more flexibility in guidance compared to commonly used approximations of the diffusion posterior.
While our method adapts well to diverse tasks, there remain several interesting directions for future work. First, a more thorough theoretical investigation into the optimization dynamics of the proposed method and integration with existing methods for fast diffusion sampling (Pandey et al., 2024a) could help alleviate the sampling costs of our method. Second, our method is only one instantiation of our framework, leaving ample room for exploration in investigating novel variational parameterizations and refining cost functions which could further enhance the flexibility of our approach.
Lastly, in a broader sense, our work gives another example of the intricate connections between test-time adaptation of diffusion models and Bayesian inference; this view may enable future generative models to perform more complex inference tasks such as hierarchical modeling, and uncertainty quantification (Jazbec et al., 2025).
Broader Impact. Despite good sample quality in a limited sampling budget, restoration can sometimes lead to artifacts in the generated sample, which can be undesirable in some domains like medical image analysis. Moreover, since our method relies on defining a terminal cost, it can be used for the synthesis of harmful content using an ill-defined cost objective.
Acknowledgment. We thank Justus Will for additional discussions and feedback. This project was funded through support from the Chan Zuckerberg Initiative. Additionally, Stephan Mandt acknowledges funding from the National Science Foundation (NSF) through an NSF CAREER Award IIS-2047418, IIS-2007719, the NSF LEAP Center, and the Hasso Plattner Research Center at UCI. Parts of this research was supported by the Intelligence Advanced Research Projects Activity (IARPA) via Department of Interior/ Interior Business Center (DOI/IBC) contract number 140D0423C0075. The U.S. Government is authorized to reproduce and distribute reprints for Governmental purposes notwithstanding any copyright annotation thereon. Disclaimer: The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies or endorsements, either expressed or implied, of IARPA, DOI/IBC, or the U.S. Government.
@@ A.1 Simplification of the NDTM Objective for DDIM
In the case of DDIM (Song et al., 2021a), the diffusion posterior is parameterized as (Eqn. 12 in Song et al. (2021a)),
@@ A.2 Continuous-Time Diffusion Trajectory Matching
By (Song et al., 2021b, Theorem 1 in Appendix A) (which follows from an application of Girsanov’s Theorem), the transient cost reads:
@@ B.1.1 Inverse Problems
B.1.2 Style Guidance with Stable Diffusion
In the context of text-to-image models like Stable Diffusion, the goal of style guidance is to generate a sample that simultaneously adheres well to a given text prompt and style features from a reference image. More specifically, given a reference style image ${\bm{r}}$, a pretrained feature extractor (like CLIP) ${\mathcal{F}}$, a pretrained decoder ${\mathcal{D}}$ such that $\hat{{\mathbf{x}}}_{0}^{t}={\mathcal{D}}(\hat{{\mathbf{z}}}_{0}^{t})$, we define $\Phi(\hat{{\mathbf{x}}}_{0}^{t})$ as,
$\Phi(\hat{{\mathbf{x}}}_{0}^{t})=\|{\mathcal{G}}({\mathcal{F}}({\bm{r}}))-{\mathcal{G}}({\mathcal{F}}(\hat{{\mathbf{x}}}_{0}^{t}))\|_{F}^{2}$ (44)
where ${\mathcal{G}}$ denotes the Gram-matrix operation and .$\|.\|_{F}$ denotes the Frobenius norm. In this formulation, the pretrained text-to-image diffusion model works as a generative prior, $p({\mathbf{x}}|{\mathbf{t}})$, where ${\mathbf{t}}$ is a text prompt embedding, and the goal is to generate samples from the posterior $p({\mathbf{x}}|{\mathbf{r}},{\mathbf{t}})$
B.2 Inverse Problems - Task Specific Hyperparameters
Here, we provide a detailed overview of different hyperparameters for the baselines considered in this work for the inverse problem tasks. We optimize all baselines and our method for the best sample perceptual quality. We use the official code implementation for RED-Diff (Mardani et al., 2024) at https://github.com/NVlabs/RED-diff, https://github.com/mandt-lab/c-pigdm, and https://github.com/sun-umn/DMPlug for running all competing baselines.
@@ B.2 Inverse Problems - Task Specific Hyperparameters
We adopt the DPS parameters from Mardani et al. (2024). More specifically, we fix the number of diffusion steps to 1000 using the DDIM sampler. We set $\eta=0.5$ for all tasks. Following Chung et al. (2022a), we set,
@@
Table 6 illustrates different hyperparameters for DPS on all tasks for the FFHQ and ImageNet datasets.
@@
Following Kawar et al. (2022), we fix $\eta=0.85$, $\eta_{b}=1.0$, and use the DDIM sampler with the number of diffusion steps set to 20 across all linear inverse problems.
@@
We set the number of diffusion steps to 20 for all tasks. It is also common to contract the reverse diffusion sampling for better sample quality by initializing the noisy state as proposed in Chung et al. (2022b). We denote the start time as $\tau$. We re-tune C-$\Pi$GDM for the best perceptual quality for all linear inverse problems. Table 6 illustrates different hyperparameters for linear inverse problems. We find that C-$\Pi$GDM fails to recover plausible images for the random inpainting task after numerous tuning attempts.
Table 5: DPS hyperparameters for different tasks
@@
Table 6: C-$\Pi$GDM hyperparameters used for different tasks. We find that C-$\Pi$GDM fails to recover plausible images for the random inpainting task after numerous tuning attempts.
@@
Task N $\gamma$ $\eta$ $\tau$ $w_{T}$ $w_{\text{s}}$ $w_{\text{c}}$
@@
B.2.4 RED-Diff (Mardani et al., 2024)
We set $\sigma_{0}=0$ with a linear weighting schedule and $lr=0.5$, $\lambda=0.25$, and perform 50 diffusion steps using the DDIM sampler for all tasks across the FFHQ and ImageNet datasets. We highlight different hyperparameters in Table 7.
@@
Task N $\gamma$ $\eta$ $\tau$ $w_{T}$ $w_{\text{s}}$ $w_{\text{c}}$ N $\gamma$ $\eta$ $\tau$ $w_{T}$ $w_{\text{s}}$ $w_{\text{c}}$
@@
B.2.5 MPGD (He et al., 2024)
Following the optimal settings in He et al. (2024), we use the setting MPGD w/o proj using the DDIM sampler with 100 diffusion steps and guidance scale set to 5.0 for all tasks.
B.2.6 NDTM (Ours)
We use the Adam optimizer (Kingma & Ba, 2017) with default hyperparameters, fixing the learning rate to 0.01 for updating the control ${\mathbf{u}}_{t}$ across all tasks and fixing the kernel learning rate in the BID task to 0.01. We refer to the loss weighting scheme in Eq. 16 as "DDIM weighting". Moreover, we use linear decay for the learning rate. We perform 50 diffusion steps using the DDIM sampler across all datasets and tasks. We tune the guidance weight $\gamma$, the number of optimization steps N, loss weighting ($w_{T}$, $w_{\text{s}}$, $w_{\text{c}}$), DDIM $\eta$ and the truncation time $\tau$ (Chung et al., 2022b) for best performance across different tasks. All these hyperparameters are listed in Table 9.
B.2.7 RB-Modulation (Rout et al., 2025)
Since RB-Modulation is a special case of NDTM with $\gamma=1.0$ and $w_{s}=w_{c}=0$, we re-run NDTM for different tasks with these settings, keeping all other hyperparameters fixed to report results for RB-Modulation.
B.3 Style Guidance Hyperparameters
B.3.1 MPGD (He et al., 2024)
Following He et al. (2024), we use the DDIM sampler with 100 steps and $\eta=1.0$ without the time reversal (Lugmayr et al., 2022) trick for fair comparisons. We set $\rho=17.5$ and the classifier-free guidance scale to 7.5
B.3.2 FreeDOM (Yu et al., 2023)
Following Yu et al. (2023), we use the DDIM sampler with 100 steps and $\eta=1.0$ without the time reversal trick. We use the classifier-free guidance scale of 5.0
B.3.3 NDTM
For NDTM, we use the DDIM sampler with 50 steps and $\eta=1.0$. We set the control learning rate to 0.002 with the Adam optimizer. The loss weightings are set to $w_{T}=1.0,w_{c}=0,w_{s}=0$ with $\gamma=4.0,N=2$ and a classifier-free guidance scale of 5.0